OpenStudy (anonymous):

simplify: x^2+2x+1/ 2x+2

hartnn (hartnn):

hi :) do you know how to factor x^2+2x+1 ?

hartnn (hartnn):

2 numbers with sum = 2 , but product = 1 are ??

OpenStudy (anonymous):

@hartnn I don't understand ? can we go through this step by step

hartnn (hartnn):

thats the first step ! to factor \(ax^2 +bx+c \), we first find 2 numbers with sum = +b and product = a*c so for x^2+2x+1 we need to find 2 numbers with sum = +2 and product = +1 which are 2 such numbers ?

OpenStudy (anonymous):

@hartnn x+1 is the only thing i can think of

hartnn (hartnn):

how about 1 and 1 1+1 = 2 and 1*1 = 1

hartnn (hartnn):

so we split 2x as x+x \(\Large x^2+2x+1 = x^2+x+x+1\) we can factor out x from 1st 2 terms \(\large x (x+1)+1(x+1)\) now we can factor out x+1 \(\large (x+1)(x+1)\) from the denominator 2x+2, we can factor out 2 \(\large 2x+2 = 2 (x+1)\) so we have \(\large \dfrac{(x+1)\cancel {(x+1)}}{2\cancel{(x+1)}} = \dfrac{x+1}{2}\) thats it! look at the steps, and ask if you have doubts in any step.